OpenStudy (anonymous):

Which sets of numbers are closed under multiplication? Choose all answers that are correct. A. whole numbers B. natural numbers C. negative integers D. integers

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@iGreen

OpenStudy (solomonzelman):

what it means to be closed under a multiplication is that if you multiply, let set even number times an even number you get an even number: \(\large\color{royalblue}{ {\rm even} ~\times~{\rm even}~=~\color{blueviolet}{{\rm even}} }\) (hence \(\normalsize\color{royalblue}{\color{blue}{{\rm even}} }\) is closed under multiplication)

OpenStudy (solomonzelman):

but, a counterexample, (saying that not every set is closed under multiplication) would be: \(\large\color{royalblue}{ ({\rm prime~number}) ~\times~({\rm prime~number})~\ne~\color{blueviolet}{{\rm {\rm prime~number}}} }\) because, \(\large\color{royalblue}{ ({\rm prime~number}) ~\times~({\rm prime~number})~=~\color{blueviolet}{{\rm {\rm composite~number}}} }\)

OpenStudy (solomonzelman):

that means that prime numbers are NOT closed under multiplication

OpenStudy (anonymous):

C and D?

OpenStudy (solomonzelman):

lets retry this

OpenStudy (anonymous):

im runing out of time for this test im doing please hurry :'(

OpenStudy (solomonzelman):

Tell me if each statement is (always) true or not: \(\large\color{teal}{ ({\rm whole~number}) ~\times~({\rm whole~ number})~=~\color{blueviolet}{{\rm whole~number}} }\) \(\large\color{teal}{ ({\rm natural~number}) ~\times~({\rm natural~ number})~=~\color{blueviolet}{{\rm natural~number}} }\) \(\large\color{teal}{ ({\rm negative~number}) ~\times~({\rm negative ~ number})~=~\color{blueviolet}{{\rm natural~number}} }\) \(\large\color{teal}{ ({\rm integer}) ~\times~({\rm integer})~=~\color{blueviolet}{{\rm integer }} }\)

OpenStudy (anonymous):

so its a and b?

OpenStudy (solomonzelman):

yes, but there is one more correct answer choice

OpenStudy (anonymous):

d?

OpenStudy (solomonzelman):

yup

OpenStudy (solomonzelman):

A, B & D.

OpenStudy (anonymous):

ok thank u!! :3

OpenStudy (solomonzelman):

yw

OpenStudy (solomonzelman):

messed up on the 3rd statement, but you got the idea... yw